OpenStudy (anonymous):

I WILL GIVE A MEDAL AND FAN WHOEVER ANSWERS THE QUESTION!!! [7.04] Choose the correct simplification of X^2/X^7 x5 x9 1/X^5 X^7 Y^12/ X^4 Y^9 WHEN I PUT / THAT MEANS ITS A FRACTION

OpenStudy (oaktree):

Think of it like this:\[\frac{ x^2 }{ x^7 } = x^2 \times x^{-7}\]

OpenStudy (oaktree):

So you add the exponents, and you get \[x^{2-7} = x^{-5} = \frac{ 1 }{ x^5 }\]

OpenStudy (oaktree):

And what the heck is up with the fourth answer anyway?

OpenStudy (anonymous):

lol thats what im saying, i knew for a fact that the last one wasnt it @OakTree